OpenStudy (anonymous):

For 2x - 3y = -7, when x = 4, What is the value of y? Please help

OpenStudy (anonymous):

2*4-3y=-7 8-3y=-7 -3y=-15 y=5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome